Can anyone help with the following problem?
What is the slope of the tangent line to the curve y = 2x^4 + x^3 + 6x at x = 1

Answer by nerdybill(7384)

What is the slope of the tangent line to the curve y = 2x^4 + x^3 + 6x at x = 1
y = 2x^4 + x^3 + 6x
y' = 8x^3 + 3x^2 + 6
y'(1) = 8(1)^3 + 3(1)^2 + 6
y'(1) = 8 + 3 + 6
y'(1) = 17 (slope when x=1)